A test charge of +2 μC is placed halfway between a charge of +6 μC and another of +4 μC separated by

10 cm.

(a) What is the magnitude of the force on the test charge?

(b) What is the direction of this force (away from or toward the +6 μC charge)?


1
Expert's answer
2021-01-07T11:00:41-0500

The force acting on a test charge is a sum of forces due to every charge. Let's calculate them separately :

  1. "F_{2; 6} = \frac{1}{4\pi\epsilon_0} \frac{q_1q_2}{r^2}=\frac{9\cdot10^9 \cdot 2\cdot10^{-6}\cdot6\cdot10^{-6}}{5^2\cdot10^{-4}}=43.2 N", both charges are positive so this force repulses the test charge from the "6 \mu C" charge.
  2. "F_{2;4} = \frac{1}{4\pi\epsilon_0} \frac{q_2 q_3}{r^2} = \frac{9\cdot10^9\cdot2\cdot10^{-6}\cdot4\cdot10^{-6}}{5^2 \cdot 10^{-4}}=28.8N", both charges are positive so this force repulses the test charge from the "4 \mu C" charge and thus is directed towards the "6\mu C" charge.


Therefore the resulting force is "F=F_{2;6}-F_{2;4}=14.4N" and is directed from the "6 \mu C" charge.